# Exponential Convergence of (Stochastic) Gradient Descent for Separable Logistic Regression

### v1

@@ 1 Introduction
We also analyze stochastic gradient descent in the separable logistic regression setting. Inspired by the techniques of Zhang et al. (2025b), we establish exponential convergence of vanilla SGD for separable logistic regression under increasing adaptive step sizes. To the best of our knowledge, this is the first result showing exponential convergence for SGD in this setting without resorting to line search or specialized adaptive procedures. Recent work by Vaswani and Babanezhad (2025) claims exponential convergence of SGD using an Armijo line search. However, the proof as currently presented is not correct due to improper conditioning on future randomness. In contrast, our results show that standard SGD with increasing (and potentially large) step sizes attains comparable guarantees over a broader tolerance regime, without requiring any line search and without knowledge of the final tolerance level. Moreover, by carefully conditioning on the filtration, we provide a corrected analysis that also resolves the issue in the SGD proof of Vaswani and Babanezhad (2025). Taken together, our results show that instability is not a prerequisite for acceleration. Carefully structured yet simple step-size growth suffices to achieve exponential convergence for both gradient descent and stochastic gradient descent in separable logistic regression.
@@ 2 Related Works
Classical analyses of gradient descent (GD) for smooth convex loss functions typically employ a fixed step-size $\eta$ as inverse of the global smoothness constant. This choice guarantees monotonic descent of the loss but leads only to sub-linear convergence rate of $\mathcal{O}(1/T)$, with $T$ as the horizon. Obtaining linear convergence rate generally requires the loss $\mathcal{L}(\cdot)$ to be strongly convex. Line-search (backtracking) methods (Vaswani and Babanezhad, 2025) adaptively tune the step size to guarantee monotonic descent, but they typically introduce additional per-iteration computational overhead. Moreover, the current SGD analysis in Vaswani and Babanezhad (2025) contains a minor flaw. To adapt to the local geometry, Polyak scheme (Polyak, 1969) chooses variable step-sizes, $\eta_{t}\coloneq\frac{\mathcal{L}(\mathbf{w}_{t})-\mathcal{L}^{\star}}{\|\nabla\mathcal{L}(\mathbf{w}_{t})\|^{2}}$, which requires the knowledge of $\mathcal{L}^{\star}$ and only results in $\mathcal{O}(1/T)$ convergence rate for smooth convex functions (He et al., 2025). Logistic loss, a smooth convex function widely used in classification, has motivated a growing body of work aimed at providing rigorous theoretical justification for its empirical success. More recent work has demonstrated that exploiting its local smoothness properties can substantially accelerate convergence. In particular, for logistic regression under linearly separable data, Axiotis and Sviridenko (2023) result in near exponential convergence (after a transient time) by employing $\eta_{t}\leq\min\left\{\frac{1}{2\mu\mathcal{L}(\mathbf{w}_{t})},\ \frac{1}{\bar{\gamma}\|\nabla\mathcal{L}(\mathbf{w}_{t})\|}\right\}$, where $\mu$ and $\bar{\gamma}$ are the constants associated with the local geometric properties of second-order robustness and multiplicative smoothness defined by Axiotis and Sviridenko (2023). A related but distinct line of research studies adaptive first-order methods that automatically tune step-sizes without explicit curvature information and obtain exponential convergence rate (Malitsky and Mishchenko, 2020) but with a dependence on the local strong convexity parameter. As large step-sizes for logistic loss with linearly separable data often outperform in practice, recent studies like Wu et al. (2024) have shown improved convergence rates of the order $\mathcal{O}(1/(\eta T))$, where $\eta=\mathcal{O}(T)$, after an initial unstable loss-oscillation phase lasting for $\mathcal{O}(T/2)$ iterations. Such an analysis was extended showing improved convergence rate with adaptive step-sizes by Zhang et al. (2025a). To the best of our knowledge, such exponential convergence rates for GD with large and non-adaptive step-sizes are not known. Moreover, these proof techniques do not extend straightforwardly to stochastic gradient descent (SGD). Wu et al. (2024) only discuss the initial unstable bound in SGD for logistic regression and give high probability bounds on the average loss as $O\left(\eta/T\right)$. The literature related to step-sizes adapted to local smoothness for SGD is also relatively sparse.
@@ 3.4 Analysis with Block Adaptive SGD
Why conditioning on Future Events in Vaswani and Babanezhad (2025) is invalid:
In the proof of Vaswani and Babanezhad (2025), the event
$G:=\Big\{\inf_{t\in[T]}f(\theta_{t})\leq\varepsilon\Big\},\qquad G^{c}:=\Big\{\inf_{t\in[T]}f(\theta_{t})\geq\varepsilon\Big\}$
is introduced, and the argument proceeds “conditioned on $G^{c}$”. We explain why this conditioning is not compatible with the per-iteration expectation calculations used in the proof. Let $(I_{t})$ be the sampled indices and $\mathcal{F}_{t}=\sigma(I_{0},\dots,I_{t-1})$ the natural filtration. For $t<T$, the event $G^{c}$ depends on future iterates $\theta_{s}$ for $s\geq t$ and hence on future indices $(I_{t},I_{t+1},\dots,I_{T-1})$, so in general $G^{c}\notin\mathcal{F}_{t}.$ Therefore conditioning on $G^{c}$ amounts to conditioning on future randomness. While $I_{t}$ is uniform conditional on $\mathcal{F}_{t}$, its law changes under conditioning on $G^{c}$. Indeed, by Bayes’ rule,
$\mathbb{P}(I_{t}=i\mid\mathcal{F}_{t},G^{c})=\frac{\mathbb{P}(G^{c}\mid\mathcal{F}_{t},I_{t}=i)}{\sum_{j=1}^{n}\mathbb{P}(G^{c}\mid\mathcal{F}_{t},I_{t}=j)},$
which is uniform only if the probabilities $\mathbb{P}(G^{c}\mid\mathcal{F}_{t},I_{t}=i)$ are equal across $i$, which need not hold. Intuitively, sampling a large-loss datapoint makes it more likely that the loss will drop below $\varepsilon$, thereby reducing the probability of $G^{c}$; hence conditioning on $G^{c}$ biases the sampling toward indices that slow progress. Consequently, per-step identities such as
$\mathbb{E}[f_{t}(\theta_{t})\mid\mathcal{F}_{t},G^{c}]=f(\theta_{t})$
or uniform drift bounds cannot be justified under this conditioning. The appearance of a telescoping inequality implying a negative expectation for a nonnegative quantity is precisely a symptom of this issue: the drift bound does not hold uniformly under conditioning on the future-dependent event $G^{c}$.
A correct approach replaces conditioning on $G^{c}$ with conditioning on stopping-time events (e.g. $\{\tau_{\varepsilon}>t\}\in\mathcal{F}_{t}$), under which the sampling distribution remains uniform and per-step expectations can be handled rigorously.
@@ 4 Experimental Results
Figure 4: Average loss values with SGD (10) for logistic regression over an MNIST subset for different pair of labels: (a) Label ‘4’ vs ‘9’. (b) Label ‘0’ vs ‘8’. (c) Label ‘2’ vs ‘6’. (d) Label ‘1’ vs ‘5’.
@@ A.1 Analysis with Gradient Descent
$c\;=\;\left(\frac{3\gamma^{2}}{4}\right)^{\frac{1}{3}},\qquad C\;=\;\frac{\left(\frac{9\gamma^{2}}{2}\right)^{\frac{2}{3}}}{S_{0}^{2}}.$
@@ A.2 Analysis with Stochastic Gradient Descent
First note $\{t<\tau\}\in\mathcal{F}_{t}$ because $\{\tau\leq t\}\in\mathcal{F}_{t}$ and $\{t<\tau\}=\{\tau\leq t\}^{c}$. Fix any $A\in\mathcal{F}_{t}$. Since $i_{t}$ is independent of $\mathcal{F}_{t}$ by (A3) and uniform on $\{1,\dots,n\}$,

### v2

# arXiv:2602.18946v2[cs.LG] 27 Feb 2026

@@ 1 Introduction
We also analyze stochastic gradient descent in the separable logistic regression setting. Inspired by the techniques of Zhang et al. (2025b), we establish exponential convergence of vanilla SGD for separable logistic regression under increasing adaptive step sizes. To the best of our knowledge, this is the first result showing exponential convergence for SGD in this setting without resorting to line search or specialized adaptive procedures. Recent work by Vaswani and Babanezhad (2025) establishes exponential convergence of SGD using an Armijo line search. However, the proof in the published ICML version contains a technical issue arising from improper conditioning on future randomness. A corrected analysis addressing this issue appeared in a subsequent revision released independently and contemporaneously with this work. In contrast, our results show that standard SGD with increasing (and potentially large) step sizes attains comparable guarantees over a broader tolerance regime, without requiring any line search and without knowledge of the final tolerance level. Moreover, by conditioning on the hitting time, which is adapted to the filtration, we obtain a corrected analysis that avoids any dependence on future randomness and relies solely on events measurable with respect to the filtration. Taken together, our results show that instability is not a prerequisite for acceleration. Carefully structured yet simple step-size growth suffices to achieve exponential convergence for both gradient descent and stochastic gradient descent in separable logistic regression.
@@ 2 Related Works
Classical analyses of gradient descent (GD) for smooth convex loss functions typically employ a fixed step-size $\eta$ as inverse of the global smoothness constant. This choice guarantees monotonic descent of the loss but leads only to sub-linear convergence rate of $\mathcal{O}(1/T)$, with $T$ as the horizon. Obtaining linear convergence rate generally requires the loss $\mathcal{L}(\cdot)$ to be strongly convex. Line-search (backtracking) methods (Vaswani and Babanezhad, 2025) adaptively tune the step size to guarantee monotonic descent, but they typically introduce additional per-iteration computational overhead. Moreover, the current SGD analysis in Vaswani and Babanezhad (2025) contains a minor flaw in the published ICML version which has been now fixed in a revised version contemporaneously and independently. To adapt to the local geometry, Polyak scheme (Polyak, 1969) chooses variable step-sizes, $\eta_{t}\coloneq\frac{\mathcal{L}(\mathbf{w}_{t})-\mathcal{L}^{\star}}{\|\nabla\mathcal{L}(\mathbf{w}_{t})\|^{2}}$, which requires the knowledge of $\mathcal{L}^{\star}$ and only results in $\mathcal{O}(1/T)$ convergence rate for smooth convex functions (He et al., 2025). Logistic loss, a smooth convex function widely used in classification, has motivated a growing body of work aimed at providing rigorous theoretical justification for its empirical success. More recent work has demonstrated that exploiting its local smoothness properties can substantially accelerate convergence. In particular, for logistic regression under linearly separable data, Axiotis and Sviridenko (2023) result in near exponential convergence (after a transient time) by employing $\eta_{t}\leq\min\left\{\frac{1}{2\mu\mathcal{L}(\mathbf{w}_{t})},\ \frac{1}{\bar{\gamma}\|\nabla\mathcal{L}(\mathbf{w}_{t})\|}\right\}$, where $\mu$ and $\bar{\gamma}$ are the constants associated with the local geometric properties of second-order robustness and multiplicative smoothness defined by Axiotis and Sviridenko (2023). A related but distinct line of research studies adaptive first-order methods that automatically tune step-sizes without explicit curvature information and obtain exponential convergence rate (Malitsky and Mishchenko, 2020) but with a dependence on the local strong convexity parameter. As large step-sizes for logistic loss with linearly separable data often outperform in practice, recent studies like Wu et al. (2024) have shown improved convergence rates of the order $\mathcal{O}(1/(\eta T))$, where $\eta=\mathcal{O}(T)$, after an initial unstable loss-oscillation phase lasting for $\mathcal{O}(T/2)$ iterations. Such an analysis was extended showing improved convergence rate with adaptive step-sizes by Zhang et al. (2025a). To the best of our knowledge, such exponential convergence rates for GD with large and non-adaptive step-sizes are not known. Moreover, these proof techniques do not extend straightforwardly to stochastic gradient descent (SGD). Wu et al. (2024) only discuss the initial unstable bound in SGD for logistic regression and give high probability bounds on the average loss as $O\left(\eta/T\right)$. The literature related to step-sizes adapted to local smoothness for SGD is also relatively sparse.
@@ 4 Experimental Results
(a) Label ‘4’ vs ‘9’.
(b) Label ‘0’ vs ‘8’.
(c) Label ‘2’ vs ‘6’.
(d) Label ‘1’ vs ‘5’.
Figure 4: Average loss values with SGD (10) for logistic regression over an MNIST subset for different pair of labels
@@ A.1 Analysis with Gradient Descent
$c\;=\;\left(\frac{3\gamma^{2}}{4}\right)^{\frac{1}{3}},\qquad C\;=\;\frac{\left(\frac{9\gamma^{2}}{2}\right)^{\frac{2}{3}}}{exp(-3+3\ln^{3}(S_{0})+\frac{3\gamma^{2}\ln(S_{0})}{2}-3\sqrt{2}\ln^{2}(S_{0})-\frac{1}{\sqrt{2}})^{\frac{1}{3}}}.$
@@ A.2 Analysis with Stochastic Gradient Descent
First note $\{t<\tau\}\in\mathcal{F}_{t}$ because $\{\tau\leq t\}\in\mathcal{F}_{t}$ and $\{t<\tau\}=\{\tau\leq t\}^{c}$. Fix any $A\in\mathcal{F}_{t}$. Since $i_{t}$ is independent of $\mathcal{F}_{t}$ and uniform on $\{1,\dots,n\}$,